# Finding outliers in Gaussian model-based clustering

### v5

@@ Abstract
Clustering, or unsupervised classification, is a task often plagued by outliers. Yet there is a paucity of work on handling outliers in clustering. Outlier identification algorithms tend to fall into three broad categories: outlier inclusion, outlier trimming, and post hoc outlier identification methods, with the former two often requiring pre-specification of the number of outliers. The fact that sample Mahalanobis distance is beta-distributed is used to derive an approximate distribution for the log-likelihoods of subset finite Gaussian mixture models. An algorithm is then proposed that removes the least plausible points according to the subset log-likelihoods, which are deemed outliers, until the subset log-likelihoods adhere to the reference distribution. This results in a trimming method, called OCLUST, that inherently estimates the number of outliers.
@@ 1 Introduction
In general, outliers, particularly gross ones, can significantly impact the parameter estimates. For example, when fitting a mixture model, outliers may draw component means towards them and inflate component covariance matrices. In the case of our simulated example, the originally spherical clusters are elongated (Figure 1). It is thus beneficial to either remove outliers or reduce their effect by accounting for them. The proposed OCLUST algorithm is designed to handle both mild and gross outliers.
@@ 2.2 General Unsupervised Outlier Detection Methods
Some traditional approaches to outlier analysis identify points with outlier scores greater than a certain threshold. For example, we can generate a score for each data point using a measure of central tendency $a$ and a measure of variability $b$. Points are considered outliers if their scores are larger than some threshold $c$, i.e., $\mathinner{\!\left\lvert\frac{x-a}{b}\right\rvert}>c$ or $x\notin(a-bc,a+bc)$. Examples include the three sigma rule ($c=3$) and the median absolute difference thresholding strategy (Hampel, 1974). Yang et al. (2019) make this method robust by applying the thresholding procedure twice, while Buzzi-Ferraris and Manenti (2011) remove outliers one-by-one by iteratively applying a clever mean threshold.
@@ 3.1 Distribution of Subset Log-Likelihoods
In this section, the distribution of Mahalanobis squared distance is used to derive the distribution of subset log-likelihoods, which forms the basis for the proposed OCLUST algorithm. We consider a subset log-likelihood to be the log-likelihood of a model fitted with $n-1$ of the data points. Formally, if we denote our complete dataset as $\mathcal{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$, then we can define the $j$th subset as the complete dataset with the $j$th point removed, $\mathcal{X}\setminus\mathbf{x}_{j}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{j-1},\mathbf{x}_{j+1},\dots,\mathbf{x}_{n}\}$. There are $n$ of these such subsets.
@@
which we will denote as $\ell_{\mathcal{X}}$ for simplicity. From (2), the log-likelihood of the model can be expressed as the sum of the logarithm of each point’s probability density. Points near the tails — corresponding to probability density values close to zero — will have a smaller contribution to the log-likelihood than points with higher probability density near the centre of the distribution. Intuitively, the larger the difference in log-likelihood between the subset and full models, the more outlying the point is, i.e., the model improved the most in its absence.
@@
In practice, Assumption 1 may be relaxed. For more information on the effect of cluster separation on the density, see Appendix A.
@@
A proof of Lemma 1 may be found in Appendix B. We will maintain Assumption 1 throughout this paper. The quantity $l_{\mathcal{X}}$ denotes the complete-data log-likelihood for the entire dataset $\mathcal{X}$, and we use $l_{\mathcal{X}\setminus\mathbf{x}_{j}}$ to denote the complete-data log-likelihood for the $j$th subset $\mathcal{X}\setminus\mathbf{x}_{j}$. Finally, we define the variable $Y_{j}=l_{\mathcal{X}\setminus\mathbf{x}_{j}}-l_{\mathcal{X}}$ as the difference between the $j$th subset complete-data log-likelihood and the complete-data log-likelihood for the entire dataset. This leads to our main result, which follows as Theorem 1.
@@
[…]
@@
where $n_{g}=\sum_{i=1}^{n}z_{ig}$ is the number of observations in the $g$th cluster, $\hat{\pi}_{g}$ is the sample proportion, $\bar{\mathbf{x}}_{g}$ is the sample mean, and $\mathbf{S}_{g}$ is the sample covariance for the $g$th cluster considering all observations in the entire dataset $\mathcal{X}$. Now consider $\hat{\pi}_{g\setminus j}$, $\bar{\mathbf{x}}_{g\setminus j}$ and $\mathbf{S}_{g\setminus j}$, the sample proportion, sample mean, and sample covariance matrix, respectively, for the $g$th cluster considering only observations in the $j$th subset $\mathcal{X}\setminus\mathbf{x}_{j}$. We will require $\hat{\pi}_{g\setminus j}\approx\hat{\pi}_{g}$, $\bar{\mathbf{x}}_{g\setminus j}\approx\bar{\mathbf{x}}_{g}$ and $\mathbf{S}_{g\setminus j}\approx\mathbf{S}_{g}$ for all $j$. Because we have $\hat{\pi}_{g\setminus j}\rightarrow\hat{\pi}_{g}$ $\bar{\mathbf{x}}_{g\setminus j}\rightarrow\bar{\mathbf{x}}_{g}$ and $\mathbf{S}_{g\setminus j}\rightarrow\mathbf{S}_{g}$ as $n_{g}\rightarrow\infty$ (proof in Appendix B.2), we need the following assumption.
@@
where $t_{j}=(\mathbf{x}_{j}-\bar{\mathbf{x}}_{h})^{\prime}\mathbf{S}_{h}^{-1}(\mathbf{x}_{j}-\bar{\mathbf{x}}_{h})$ is the Mahalanobis distance for point $\mathbf{x}_{j}$ using the sample parameter estimates in (6) when $z_{jh}=1$.
@@
[…]
@@
Because $W_{j}$ is beta distributed, it has density of the from (3). A transformation of variables allows the density of $Y_{j}\mid(z_{jh}=1)$ to be written
@@ 3.2 OCLUST Algorithm
Consider $y_{1},\ldots,y_{n}$ to be realizations of the random variable $Y$, which has density given by (11). We propose testing the adherence of $y_{1},\ldots,y_{n}$ to the reference distribution in (11) as a way to test for the presence of outliers. In other words, if $y_{1},\ldots,y_{n}$ does not have a beta mixture distribution, then we assume outliers are present in the model. Because ${\ell}_{\mathcal{X}}$ converges to $l_{\mathcal{X}}$ and acts as a good approximation, we will use $\ell_{\mathcal{X}}$. This is important because we will need $\ell_{\mathcal{X}\setminus\mathbf{x}_{j}}$ for outlier identification and, additionally, it is outputted by many existing clustering algorithms. The algorithm described below uses the log-likelihood and parameter estimates calculated using the expectation-maximization algorithm (EM; Dempster et al., 1977) for Gaussian model-based clustering; however, other methods may be used for parameter estimation. The proposed algorithm is called OCLUST. It both identifies likely outliers and estimates the proportion of outliers within the dataset. When $\mathbf{X}$ is distributed according to a Gaussian mixture model, then ${Y}$ has the probability density in (11). If $y_{1},\ldots,y_{n}$ does not come from (11), then $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ does not arise from a Gaussian mixture model (1), implying that the model is misspecified. We assume the misspecification is due to outliers. The OCLUST algorithm assumes that the model is otherwise correctly specified and when $Y$ does not follow the distribution in (11), then outliers are present. The ‘closeness’ of the distribution of $Y$ to (11) is assessed using the Kullback-Leibler (KL) divergence, estimated via relative frequencies. The algorithm (Algorithm 1) involves removing candidate outliers one-by-one until KL divergence is minimized. At each step, we obtain the log-likelihood of the entire dataset and the log-likelihoods of the $n$ subsets. We then assess the distribution of subset log-likelihoods and determine our next candidate outlier, $t$, according to Definition 1. This candidate outlier is then trimmed and the algorithm continues to the next iteration. Notably, KL divergence generally decreases as outliers are removed and the model improves. Once all outliers are removed, KL divergence increases again as points are removed from the tails. We select the number of outliers as the location of the minimum KL divergence. The OCLUST algorithm is outlined in Algorithm 1, with $n$ being the size of the dataset, $G$ the number of clusters, $B\in[0,F-1]$ the number of initially rejected outliers (see Remark 2), and $F$ the chosen upper bound (see Remark 3). An R implementation of the algorithm is available as the oclust package on CRAN (Clark and McNicholas, 2022).
@@
     clustering algorithm, e.g., the EM algorithm in the mixture R package
@@ 3.3 Computational Complexity
Clustering must be performed $n~{}+~{}1$ times for each of the $F+1$ iterations of the algorithm. In our simulation study, we use the mixture package to implement the EM algorithm, the order of which is generally $O(Gp^{3}n)$. As a result, OCLUST is $O(FGp^{3}n^{2})$. The computation time can be reduced by initializing the subset models with the parameters from the full model to vastly decrease the number of iterations required within the clustering algorithm. In addition, subset models are independent and can easily be computed in parallel. Computation times for applying OCLUST to the datasets in Section 4 are available in Appendix C. These times are similar to or somewhat faster than the mean-shift outlier detection algorithm when OCLUST is run in parallel. It is important to note that these times for OCLUST take into account the algorithm running to the upper bound $F$, which is often unnecessary. Thus, we propose an early stopping criterion below. The original version of the OCLUST algorithm involves calculating the KL divergence over a range of possible numbers of outliers. The best model is then chosen as the one which minimizes the KL divergence. Instead, consider a stopping rule for the algorithm to minimize the number of iterations required. We can stop the algorithm once the data fit well to the distribution. One way to do this is using Kuiper’s Test (Kuiper, 1960), where
@@
where $D^{+}=\max[F_{E}(y_{i})-F_{0}(y_{i})]$ and $D^{-}=\max[F_{0}(y_{i})-F_{E}(y_{i})]$, for $i\in[1,n]$, $F_{0}(y_{i})$ is the CDF of $Y$ at point $y_{i}$, and $F_{E}(y)$ is the empirical CDF at point $y_{i}$ for a sample of size $n$, i.e., $F_{E}(y_{i})=i/n$. An approximate p-value can be estimated using Monte Carlo simulation from the proposed distribution in (11). A total of $B$ datasets of size $n$ are simulated and the test statistic is calculated for each one. Then, we estimate the p-value as $(r+1)/(B+1)$, where $B$ is the number of samples generated and $r$ is the number of samples to produce a test statistic greater than or equal to $T_{0}$ (North et al., 2002). Instead of using KL divergence in the OCLUST algorithm, we calculate the approximate p-value for each iteration and stop once it is higher than a pre-specified significance level, e.g., 5% or 10%.
@@ 4.2 Simulation Study
The following simulation study tests the performance of OCLUST against the following five popular outlier detection algorithms:
@@
OCLUST is a trimming method which removes outliers from the model. In contrast, CNMix and NCM are outlier inclusion methods. CNMix treats outliers as contamination and assumes the outliers are cluster-specific and reside around the true clusters. NCM assumes that outliers are included uniformly. DBSCAN treats clusters as density-connected points and outliers as points that do not belong to clusters. Mean-shift outlier detection is a pre-processing step for a subsequent clustering algorithm. Finally, the thresholding approach considers a point an outlier if the outlier ‘score’ exceeds a specified threshold. Following Yang et al. (2019), we apply this procedure twice (2T). Calculated before clustering, we choose the outlier ‘score’ to be the Mahalanobis distance from the centre of the dataset. This measure accounts for differences in variability among dimensions. Subsequently, we cluster the remaining data using the EM algorithm in the mixture package. These methods are broadly classified into two types: those that remove outliers and then cluster (mean-shift and 2T) and those that do both simultaneously (OCLUST, CNMix, NCM, DBSCAN). The following simulation scheme closely follows that in Fränti and Yang (2018). We use the same eight two-dimensional benchmark datasets plus one 32-dimensional dataset (Fränti and Sieranoja, 2018). Sets S1–S4 investigate increasing degrees of overlap. There is a mix of 15 spherical and non-spherical Gaussian clusters with overlap ranging from 9% to 44%. Sets A1–A3 have increasing numbers of clusters, with all sets having spherical Gaussian clusters with equal cluster sizes, deviation, and overlap. The Unbalance set has three clusters densely populated with 2000 points each and five clusters sparsely populated with 100 points each. The Dim032 dataset has 16 well-separated Gaussian clusters in 32 dimensions. The datasets are summarized in Table 1 and plotted in Figure 3, with a two-dimensional projection shown for the Dim032 dataset.
@@
Figure 3: The nine benchmark datasets used herein, where red points represent added uniform noise and other colours correspond to the true clusters.
@@
Following Fränti and Yang (2018), we set the proportion of outliers at 7%. We generate uniform noise in each dimension in the range $[x_{\text{mean}}-2\times\text{range},x_{\text{mean}}+2\times\text{range}]$, where the range is the difference between the farthest observation and the mean in each dimension. We run oclust in R, fixing the upper bound $F$ from Table 1, i.e., at or about 10% of the original data size. We eliminate initial gross outliers with large Mahalanobis distances to regions of high density identified by the dbscan package (Hahsler et al., 2019). Covariance matrices for the S sets, A sets, and Unbalance set are unrestricted, allowing varying volumes, shapes, and orientations among the clusters. We restrict the Dim032 set to a variance structure with varying volumes and spherical shapes to hasten computation time. We run CNMix using the CNmixt function from the ContaminatedMixt package (Punzo et al., 2018) with default $k$-means initialization. The CNmixt function chooses the best model for variance and contamination using the BIC. We run NCM using the Mclust function in the mclust package (Scrucca et al., 2016), initializing the noise component as a random sample of points with probability 1/4. We let the package choose the best model using the Bayesian information criterion (Schwarz, 1978). We run mean-shift outlier detection in Python using Yang (n.d.)’s code. We specify the number of clusters and perform three iterations. Due to its superior performance in Fränti and Yang (2018), non-outlying points are clustered using random swap clustering in Python (Yang et al., 2021; Fränti, 2018) and run for the recommended 5000 iterations. We implement the DBSCAN algorithm with the dbscan package, setting the minimum number of points per cluster to be 20. For the thresholding approach, we calculate Mahalanobis distance from the centre of the dataset and trim those that are more than three standard deviations away from the mean. The remaining points are clustered with the mixture package. The performance of each method can be assessed using three approaches for evaluating classification accuracy — adjusted Rand index (ARI; Hubert and Arabie, 1985), normalized mutual information (NMI; Kvalseth, 1987), and centroid index (CI; Fränti et al., 2014) — and two approaches of outlier detection performance — true positive rate (TPR) and false positive rate (FPR). The ARI is a measure of agreement between two partitions, e.g., between predicted clusters and true classes. The ARI equals 1 with perfect agreement and has expected value 0 when randomly partitioned. Negative values indicate partitions with less agreement than by random chance. NMI is a measure of the amount of information shared between the true classes and the clustering solution. It is a scale from 0 to 1, with 0 corresponding to no mutual information and 1 corresponding to perfect agreement. CI is a measure of dissimilarity among cluster centroids, where $\text{CI}=0$ reflects one-to-one correspondence between the estimated cluster centres and the ground truth. We calculate ARI and NMI for cluster agreement taking into account whether points are classified into the correct group, considering outliers as one of those groups. TPR reflects the proportion of true outliers classified as such and FPR reflects the proportion of non-outliers classified as outliers. The results are shown in Table 2, with the best algorithm according to each approach bolded.
@@
For datasets with many clusters, i.e., A2 and A3, OCLUST does better than its competitors. OCLUST and mean-shift have $\text{CI}=0$ for all but one dataset. Mean-shift performs better than the Gaussian model-based approaches as the clusters become highly overlapped in S4. The model-based approaches cannot recover symmetric distributions, such as the Gaussian distribution, for each component when the data severely overlap. The OCLUST and mean-shift algorithms attain similar results on the A1, S1, S2, and S3 datasets. Mean-shift performs well, as expected, for those datasets because the clusters are spherical and well-separated. With the exception of DBSCAN, most of the algorithms have a nearly zero FPR. DBSCAN has a perfect TPR for four of the datasets, but this is also accompanied by a very large FPR, indicating that DBSCAN is considering most points to be outliers for those datasets. After discarding these results from DBSCAN, OCLUST has the largest TPR for most of the datasets. OCLUST, mean-shift, and NCM perform very well on the 32-dimensional dataset, likely due to the uniform generation of outliers. This is because, as the dimensionality increases, it is more likely for a generated outlier to be far from the clusters in any one direction. OCLUST, CNMix, MCN, and DBSCAN perform well on the Unbalance dataset. Mean-shift filters most of the Unbalance points as noise, resulting in poor classification. DBSCAN performs poorly on most datasets, often over-specifying the number of outliers and combining clusters together. The thresholding approach performs consistently for outlier identification but fails when clusters become more overlapped, likely due to remaining outliers reducing clustering performance. In Figure 4, the KL graph for the A3 dataset is plotted. It displays a typical shape, decreasing to a minimum at 446 outliers, after which KL increases. The minimum at 446 is notably different from 525, the number of outliers generated. However, this is to be expected due to the uniform generation of outliers because many of them overlap with the clusters as seen in Figure 3. All the algorithms experience the same issue in this respect.
@@
In the initial gross outlier removal for the A3 dataset, 399 outliers are removed. We reach a minimum KL divergence after 48 iterations, but continuing until $F=750$ requires a further 304 iterations. At the minimum KL divergence, we achieve an approximate p-value of 0.099 on Kuiper’s test when $B=100$. If our significance level was 0.05, we would be able to stop here and substantially save computation time.

### v6

@@ Abstract
Clustering, or unsupervised classification, is a task often plagued by outliers. Yet there is a paucity of work on handling outliers in clustering. Outlier identification algorithms tend to fall into three broad categories: outlier inclusion, outlier trimming, and post hoc outlier identification methods, with the former two often requiring pre-specification of the number of outliers. The fact that sample squared Mahalanobis distance is beta-distributed is used to derive an approximate distribution for the log-likelihoods of subset finite Gaussian mixture models. An algorithm is then proposed that removes the least plausible points according to the subset log-likelihoods, which are deemed outliers, until the subset log-likelihoods adhere to the reference distribution. This results in a trimming method, called OCLUST, that inherently estimates the number of outliers.
@@ 1 Introduction
In general, outliers, particularly gross ones, can significantly impact the parameter estimates. For example, when fitting a mixture model, outliers may draw component means towards them and inflate component covariance matrices. In the case of our simulated example, the originally-spherical clusters are elongated (Figure 1). It is thus beneficial to either remove outliers or reduce their effect by accounting for them. The proposed OCLUST algorithm is designed to handle both mild and gross outliers.
@@ 2.2 General Unsupervised Outlier Detection Methods
[…]
@@ 3.1 Distribution of Subset Log-Likelihoods
In this section, the distribution of squared Mahalanobis distance is used to derive the distribution of subset log-likelihoods, which forms the basis for the proposed OCLUST algorithm. We consider a subset log-likelihood to be the log-likelihood of a model fitted with $n-1$ of the data points. Formally, if we denote our complete dataset as $\mathcal{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$, then we can define the $j$th subset as the complete dataset with the $j$th point removed, $\mathcal{X}\setminus\mathbf{x}_{j}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{j-1},\mathbf{x}_{j+1},\dots,\mathbf{x}_{n}\}$. There are $n$ such subsets.
@@
which we will denote as $\ell_{\mathcal{X}}$ for simplicity. From (2), the log-likelihood of the model can be expressed as the sum of the logarithm of each point’s probability density. Points near the tails — corresponding to probability density values close to zero — will have a smaller contribution to the log-likelihood than points with higher probability density such as points near the centre of the distribution. Intuitively, the larger the difference in log-likelihood between the subset and full models, the more outlying the point is, i.e., the model improved the most in its absence.
@@
[…]
@@
[…]
@@
$Y_{j}\mid(z_{jh}=1)\sim f_{\text{beta}}\left(\frac{2n_{h}}{(n_{h}-1)^{2}}(y_{j}-c)~{}\bigg{|}~{}\frac{p}{2},\frac{n_{h}-p-1}{2}\right)$ (5)
@@
[…]
@@
where $t_{j}=(\mathbf{x}_{j}-\bar{\mathbf{x}}_{h})^{\prime}\mathbf{S}_{h}^{-1}(\mathbf{x}_{j}-\bar{\mathbf{x}}_{h})$ is the squared Mahalanobis distance for point $\mathbf{x}_{j}$ using the sample parameter estimates in (6) when $z_{jh}=1$.
@@
Sample squared Mahalanobis distance is distributed according to a scaled beta distribution (Gnanadesikan and Kettenring, 1972). When the data are multivariate normally distributed, i.e., $\mathbf{X}\sim\text{MVN}\left(\boldsymbol{\mu},\mathbf{\Sigma}\right)$,
$\frac{n}{(n-1)^{2}}T_{j}\sim f_{\text{beta}}\left(\frac{n}{(n-1)^{2}}t_{j}~{}\Bigg{|}~{}\frac{p}{2},\frac{n-p-1}{2}\right)$
@@
Because $W_{j}$ is beta distributed, it has density of the from (3). The change of variables allows the density of $Y_{j}\mid(z_{jh}=1)$ to be written
@@ 3.2 OCLUST Algorithm
Consider $y_{1},\ldots,y_{n}$ to be realizations of the random variable $Y$, which has density given by (11). We propose testing the adherence of $y_{1},\ldots,y_{n}$ to the reference distribution in (11) as a way to test for the presence of outliers. In other words, if $y_{1},\ldots,y_{n}$ does not have a beta mixture distribution, then we assume outliers are present in the model. Because ${\ell}_{\mathcal{X}}$ converges to $l_{\mathcal{X}}$ and acts as a good approximation, we will use $\ell_{\mathcal{X}}$. This is important because we will need $\ell_{\mathcal{X}\setminus\mathbf{x}_{j}}$ for outlier identification and, additionally, it is outputted by many existing clustering algorithms. The algorithm described below uses the log-likelihood and parameter estimates calculated using the expectation-maximization (EM) algorithm (Dempster et al., 1977) for Gaussian model-based clustering; however, other methods may be used for parameter estimation.
The proposed algorithm is called OCLUST. It both identifies likely outliers and estimates the proportion of outliers within the dataset. When $\mathbf{X}$ is distributed according to a Gaussian mixture model, then ${Y}$ has the probability density in (11). If $y_{1},\ldots,y_{n}$ does not come from (11), then $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ does not arise from a Gaussian mixture model (1), implying that the model is misspecified. We assume the misspecification is due to outliers. The OCLUST algorithm assumes that the model is otherwise correctly specified and when $Y$ does not follow the distribution in (11), then outliers are present. The ‘closeness’ of the distribution of $Y$ to (11) is assessed using the Kullback-Leibler (KL) divergence, estimated via relative frequencies. The algorithm (Algorithm 1) involves removing candidate outliers one-by-one until KL divergence is minimized. At each step, we obtain the log-likelihood of the entire dataset and the log-likelihoods of the $n$ subsets. We then assess the distribution of subset log-likelihoods and determine our next candidate outlier, $t$, according to Definition 1. This candidate outlier is then trimmed and the algorithm continues to the next iteration. Notably, KL divergence generally decreases as outliers are removed and the model improves. Once all outliers are removed, KL divergence increases again as points are removed from the tails. We select the number of outliers as the location of the minimum KL divergence.
The OCLUST algorithm is outlined in Algorithm 1, with $n$ being the size of the dataset, $G$ the number of clusters, $B\in[0,F-1]$ the number of initially rejected outliers (see Remark 2), and $F$ the chosen upper bound (see Remark 3). An R implementation of the algorithm is available as the oclust package on CRAN (Clark and McNicholas, 2022).
@@
     clustering algorithm, e.g., the EM algorithm in the mixture package
@@ 3.3 Computational Complexity
[…]
The original version of the OCLUST algorithm involves calculating the KL divergence over a range of possible numbers of outliers. The best model is then chosen as the one which minimizes the KL divergence. Instead, consider a stopping rule for the algorithm to minimize the number of iterations required. We can stop the algorithm once the data fit well to the distribution. One way to do this is using Kuiper’s Test (Kuiper, 1960), where
@@
where $D^{+}=\max[F_{E}(y_{i})-F_{0}(y_{i})]$ and $D^{-}=\max[F_{0}(y_{i})-F_{E}(y_{i})]$, for $i\in[1,n]$, $F_{0}(y_{i})$ is the CDF of $Y$ at point $y_{i}$, and $F_{E}(y)$ is the empirical CDF at point $y_{i}$ for a sample of size $n$, i.e., $F_{E}(y_{i})=i/n$. An approximate p-value can be estimated using Monte Carlo simulation from the proposed distribution in (11). A total of $m$ datasets of size $n$ are simulated and the test statistic is calculated for each one. Then, we estimate the p-value as $(r+1)/(m+1)$, where $m$ is the number of samples generated and $r$ is the number of samples to produce a test statistic greater than or equal to $T_{0}$ (North et al., 2002). Instead of using KL divergence in the OCLUST algorithm, we calculate the approximate p-value for each iteration and stop once it is higher than a pre-specified significance level, e.g., 5% or 10%.
@@ 4.2 Simulation Study
The following simulation study tests the performance of OCLUST against the following five popular algorithms:
@@
OCLUST is a trimming method which removes outliers from the model. In contrast, CNMix and NCM are outlier inclusion methods. CNMix treats outliers as contamination and assumes the outliers are cluster-specific and reside around the clusters. NCM assumes that outliers are included uniformly. DBSCAN treats clusters as density-connected points and outliers as points that do not belong to clusters. Mean-shift outlier detection is a pre-processing step for a subsequent clustering algorithm. Finally, the thresholding approach considers a point an outlier if the outlier ‘score’ exceeds a specified threshold. Following Yang et al. (2019), we apply this procedure twice (2T). Calculated before clustering, we choose the outlier ‘score’ to be the squared Mahalanobis distance from the centre of the dataset. This measure accounts for differences in variability among dimensions. Subsequently, we cluster the remaining data using the EM algorithm in the mixture package. These methods are broadly classified into two types: those that remove outliers and then cluster (mean-shift and 2T) and those that do both simultaneously (OCLUST, CNMix, NCM, DBSCAN).
The following simulation scheme closely follows that in Fränti and Yang (2018). We use the same eight two-dimensional benchmark datasets plus one 32-dimensional dataset (Fränti and Sieranoja, 2018). Sets S1–S4 investigate increasing degrees of overlap. There is a mix of 15 spherical and non-spherical Gaussian clusters with overlap ranging from 9% to 44%. Sets A1–A3 have increasing numbers of clusters, with all sets having spherical Gaussian clusters with equal cluster sizes, deviation, and overlap. The Unbalance set has three clusters densely populated with 2000 points each and five clusters sparsely populated with 100 points each. The Dim032 dataset has 16 well-separated Gaussian clusters in 32 dimensions. The datasets are summarized in Table 1 and plotted in Figure 3, with a two-dimensional projection shown for the Dim032 dataset.
@@
Figure 3: The nine benchmark datasets used herein, where red points represent added uniform noise and other colours correspond to the true classes.
Following Fränti and Yang (2018), we set the proportion of outliers at 7%. We generate uniform noise in each dimension in the range $[x_{\text{mean}}-2\times\text{range},x_{\text{mean}}+2\times\text{range}]$, where the range is the difference between the farthest observation and the mean in each dimension.
@@
We run oclust in R, fixing the upper bound $F$ from Table 1, i.e., at or about 10% of the original data size. We eliminate initial gross outliers with large squared Mahalanobis distances to regions of high density identified by the dbscan package (Hahsler et al., 2019). Covariance matrices for the S sets, A sets, and Unbalance set are unrestricted, allowing varying volumes, shapes, and orientations among the clusters. We restrict the Dim032 set to a variance structure with varying volumes and spherical shapes to hasten computation time. We run CNMix using the CNmixt function from the ContaminatedMixt package (Punzo et al., 2018) with default $k$-means initialization. The CNmixt function chooses the best model for variance and contamination using the BIC. We run NCM using the Mclust function in the mclust package (Scrucca et al., 2016), initializing the noise component as a random sample of points with probability 1/4. We let the package choose the best model using the Bayesian information criterion (Schwarz, 1978). We run mean-shift outlier detection in Python using Yang (n.d.)’s code. We specify the number of clusters and perform three iterations. Due to its superior performance in Fränti and Yang (2018), non-outlying points are clustered using random swap clustering in Python (Yang et al., 2021; Fränti, 2018) and run for the recommended 5000 iterations. We implement the DBSCAN algorithm with the dbscan package, setting the minimum number of points per cluster to be 20. For the thresholding approach, we calculate the squared Mahalanobis distance from the centre of the dataset and trim those that are more than three standard deviations away from the mean. This is performed twice, and the remaining points are clustered with the mixture package.
The performance of each method can be assessed using three approaches for evaluating classification accuracy — adjusted Rand index (ARI; Hubert and Arabie, 1985), normalized mutual information (NMI; Kvalseth, 1987), and centroid index (CI; Fränti et al., 2014) — and two approaches of outlier detection performance — true positive rate (TPR) and false positive rate (FPR). The ARI is a measure of agreement between two partitions, e.g., between clusters and true classes. The ARI equals 1 with perfect agreement and has expected value 0 when randomly partitioned. Negative values indicate partitions with less agreement than would be expected by random chance. NMI is a measure of the amount of information shared between the true classes and the clustering solution. It is a scale from 0 to 1, with 0 corresponding to no mutual information and 1 corresponding to perfect agreement. CI is a measure of dissimilarity among cluster centroids, where $\text{CI}=0$ reflects one-to-one correspondence between the estimated cluster centres and the ground truth. We calculate ARI and NMI for cluster agreement taking into account whether points are classified into the correct group, considering outliers as one of those groups. TPR reflects the proportion of true outliers classified as such and FPR reflects the proportion of non-outliers classified as outliers. The results are shown in Table 2, with the best algorithm according to each approach bolded.
@@
For datasets with many clusters, i.e., A2 and A3, OCLUST does better than its competitors. OCLUST and mean-shift have $\text{CI}=0$ for all but one dataset. Mean-shift performs better than the Gaussian model-based approaches as the clusters become highly overlapped in S4. The model-based approaches cannot recover symmetric distributions, such as the Gaussian distribution, for each component when the clusters severely overlap. The OCLUST and mean-shift algorithms attain similar results on the A1, S1, S2, and S3 datasets. Mean-shift performs well, as expected, for those datasets because the clusters are spherical and well-separated. With the exception of DBSCAN, most of the algorithms have a nearly zero FPR. DBSCAN has a perfect TPR for four of the datasets, but this is also accompanied by a very large FPR, indicating that DBSCAN is considering most points to be outliers for those datasets. After discarding these results from DBSCAN, OCLUST has the largest TPR for most of the datasets.
OCLUST, mean-shift, and NCM perform very well on the 32-dimensional dataset, likely due to the uniform generation of outliers. This is because, as the dimensionality increases, it is more likely for a generated outlier to be far from the clusters in any one direction. OCLUST, CNMix, NCM, and DBSCAN perform well on the Unbalance dataset. Mean-shift filters most of the Unbalance points as noise, resulting in poor classification. DBSCAN performs poorly on most datasets, often over-specifying the number of outliers and combining clusters together. The thresholding approach performs consistently for outlier identification but fails when clusters become more overlapped, likely due to remaining outliers reducing clustering performance.
In Figure 4, the KL graph for the A3 dataset is plotted. It displays a typical shape, decreasing to a minimum at 446 outliers, after which KL increases. The minimum at 446 is notably different from 525, the number of outliers generated. However, this is to be expected due to the uniform generation of outliers because many of them overlap with the clusters as seen in Figure 3. All the algorithms experience the same issue in this respect.
@@
In the initial gross outlier removal for the A3 dataset, 399 outliers are removed. We reach a minimum KL divergence after 48 iterations, but continuing until $F=750$ requires a further 304 iterations. At the minimum KL divergence, we achieve an approximate p-value of 0.099 on Kuiper’s test when $m=100$. If our significance level was 0.05, we would be able to stop here and substantially save computation time.